# TerDiT: Ternary Diffusion Models with Transformers

### v1

@@ 3.3 QAT-specific Model Structure Improvement
The large-activation problem brought about by the ternary linear weights can be alleviated by applying a layer norm to the output of the ternary linear layer. We add an RMS Norm (similar to LLaMA) after the ternary linear layer and obtain the activation distribution (as shown in the left part of Fig. 3). In this case, the activation values are scaled to a reasonable range after passing the normalization layer and lead to more stable training behavior. The observation also aligns with [17], where a layer normalization function is applied before the activation quantization for each quantized linear layer.
RMS Normalized AdaLN Module. We analyze the DiT model for QAT-specific model structure improvement based on the above insights. In a standard ViT transformer block, the layer norm is applied for every self-attention layer and feedforward layer. This is also the case with the self-attention layers and feedforward layers in the DiT block, which can help to properly scale the range of activations. However, the DiT block differs from traditional transformer blocks due to the presence of the AdaLN module, as introduced in Sec. 3.1. Notably, there is no layer normalization applied to this module. In the context of full-precision training, the absence of layer normalization does not have a significant impact. However, for ternary DiT networks, its absence can result in large dimension-wise scale and shift values in the adaLN (normalization) module, posing bad influences on model training. To mitigate this issue, we introduce an RMS Norm after the MLP layer of the adaLN module in each ternary DiT block:
[…]
and the final model structure of TerDiT is illustrated in Fig. 2 (A). This minor modification can result in faster convergence speed and lower training loss, leading to better quantitative and qualitative evaluation results. To better showcase the effect, the actual activation distribution with/without the RMS Norm after model training is analyzed in Sec. A.1.
3.4 Deployment Scheme
After training the DiT model, we find that there are currently no effective open-source deployment solutions for ternary networks. In this case, we deploy the trained networks with a 2-bit implementation. To be specific, we pack the ternary linear weights to int8 values (4 ternary numbers into one int8 number) with the pack_2bit_u8() function provided by [44]. During the inference process of the DiT model, we call the complementary unpack_2bit_u8() function on the fly to recover packed 2-bit numbers to floating-point values, then perform subsequent calculations. The addition of the unpacking operation will slow down the inference process, but we believe that with further research into model ternarization, more hardware support for inference speedup will become available.
4 Experiments
In this section, a set of experiments are carried out to evaluate our proposed TerDiT. We show our main evaluation results in Sec. 4.1, carry out deployment efficiency comparison in Sec. 4.2, and illustrate the effectiveness of the RMS Normalized adaLN module in Sec. 4.3. Our DiT implementation is based on the open-sourced code of Large-DiT-ImageNet [^3] https://github.com/Alpha-VLLM/LLaMA2-Accessory/tree/main/Large-DiT-ImageNet. We conduct experiments on ternary DiT models with 600M (size of DiT-XL/2) and 4.2B [^4] The provided 3B model in the Large-DiT-ImageNet repository actually has 4.2B parameters. (size of Large-DiT-4.2B) parameters respectively.
4.1 Main Evaluation Results
We provide quantitative and qualitative evaluation results of TerDiT in this subsection. To the best of our knowledge, there is no published work done on the quantization of diffusion transformer models at this point, so we mainly focus on the comparison with representative full-precision diffusion models in this subsection.
Remark on TerDiT Baselines. There is still no work studying the quantization of DiT models to our best knowledge. Other than comparison with full-precision models in this subsection, we have also established some baselines for comparison in other subsections. For the QAT baseline, we directly train the ternary DiT model without RMS Norm in the adaLN module in Sec. 4.3. To compare with existing PTQ [45] methods, we perform 4-bit weight quantization for pre-trained models on the same set of parameters as TerDiT, and find they just fail to generate viewable images (detailed in Sec. A.2).
Experiment Setup. Following the evaluation setting of the original DiT paper [6], we train 600M and 4.2B ternary DiT models on the ImageNet dataset. Due to computation resource limitations, we train and evaluate the model under 256$\times$256 resolution, but we think that the evaluation results are already quite representative. We compare TerDiT with a series of full-precision diffusion models, report FID [46], sFID [47], Inception Score, Precision, and Recall (50k generated images) following [48]. We also provide the total number of images (million) during the training stage as in [23] to offer further insights into the convergence speed of different generative models.
@@ 4.1 Main Evaluation Results
Table 1: Comparison between TerDiT and a series of full-precision diffusion models on the ImageNet $256\times 256$ label-conditional generation task. For generation with classifier-free guidance, we use cfg=1.5. As can be seen, TerDiT achieves comparable results with full-precision models.
Checkpoint Size Memory Usage Inference Time FID-50k
DiT-XL/2-G 2.6GB 3921MiB 15s 2.27
Large-DiT-4.2B-G 16GB 17771MiB 83s 2.10
@@
TerDiT-4.2B-G 1.1GB 3011MiB 97s 2.42
Table 2: Deployment efficiency comparison with cfg=1.5. TerDiT achieves a significant reduction in model size and memory usage while maintaining competitive evaluation results. Although the inference time of TerDiT is slightly slower due to the unpack operations, the inference time is expected to be significantly reduced with hardware support specifically designed for ternary models.
Training Details. We train the 600M TerDiT model on 8 A100-80G GPUs for 1750k steps with batchsize set to 256, and the 4.2B model on 16 A100-80G GPUs for 1180k steps with batchsize set to 512. We set the initial learning rate as 5e-4555The default learning rate of DiT is 1e-4. In our experiments, we increase the initial learning rate to 5e-4 following the observation in [17] that a larger learning rate is needed for the faster convergence of low-bit QAT.. After 1550k steps of training for the 600M model and 550k steps for the 4.2B model, we reduce the learning rate back to 1e-4 for more fine-grained parameter updates (ablation study on this learning rate reduction is provided in Sec. A.3).
Quantitative Results Analysis. The evaluation results are listed in Tab. 4.1. TerDiT is a QAT scheme for DiT models, so among all the full-precision models, we pay special attention to DiT-XL/2 (675M) and Large-DiT-4.2B. Without classifier-free guidance, TerDiT-4.2B achieves very similar testing results with DiT-XL/2 (with a much fewer number of training images). With classifier-free guidance (cfg=1.5), TerDiT-4.2B-G outperforms LDM-G while bringing a very slight performance degradation compared with two full-precision DiT-structured models. Besides, TerDiT-4.2B-G achieves better evaluation results than TerDiT-600M-G, implying that models with more parameters can incur smaller performance degradation after quantization.
Qualitative Results Analysis. To visually demonstrate the effectiveness of TerDiT, we also show some qualitative comparison results in Fig. 4, concerning TerDiT-4.2B, DiT-XL/2 and Large-DiT-4.2B. In terms of visual perception, there is no significant difference between the images generated by TerDiT and those generated by the full-precision models.
@@
Figure 4: Qualitative results analysis. We compare images generated by DiT-XL/2 (A), Large-DiT-4.2B (B), and TerDiT-4.2B (C). We use class labels [355, 980, 328, 33, 928, 862, 330, 944] and cfg=4. TerDiT-4.2B can generate images of the same quality as the other two full-precision DiT models.
4.2 Deployment Efficiency Comparison
@@
The improvement in deployment efficiency is the motivation of our proposed TerDiT scheme. In this subsection, we provide a comparison between TerDiT-600M/4.2B, DiT-XL/2, and Large-DiT-4.2B to discuss the actual deployment efficiency TerDiT can bring about. Tab. 4.1 shows the checkpoint sizes of the four DiT models. We also record the memory usage and inference time of the total diffusion sample loop (step=250) on a single A100-80G GPU.
@@
From the table, we can see that TerDiT greatly reduces checkpoint size and memory usage. The checkpoint size and memory usage of the 4.2B ternary DiT model are significantly smaller than those of Large-DiT-4.2B, even smaller than DiT-XL/2. This brings significant advantages to deploying the model on end devices (e.g., mobile phones). We observe a slower inference speed due to the required unpack operations, but we believe that the computational advantage of ternary weight networks will be showcased through future hardware support.
4.3 Discussion on the RMS Normalized AdaLN Module
Experiment Setup. We train ternary DiT models with 600M and 4.2B parameters on the ImageNet [24] dataset in 256$\times$256 resolution. For each parameter size, we train two models, one with RMS Norm in the adaLN module and one without (our baseline). We record the loss curves during training and measure the FID-50k score (cfg=1.5) every 100k training steps.
Training Details. For a fair comparison, we train all the ternary DiT models on 8 A100-80G GPUs with batchsize set to 256. The learning rate is set to 5e-4 throughout the training process.
Results Analysis. The training loss [^6] The actual training process of diffusion models is not as ‘smooth’ as one might assume. To better illustrate the training dynamics, we employ exponential smoothing (with a smoothing factor of 0.995) during visualization. and evaluation scores are shown in Fig. 5 and Fig. 6 respectively. As can be seen, training with the RMS Normalized adaLN module will lead to faster convergence speed and lower FID scores. Another observation is that models with more parameters tend to achieve faster and better training compared to models with fewer parameters. This also to some extent reflects the scaling law of the ternary DiT model. Qualitative comparison results are shown in Sec. A.4.
@@ 4.3 Discussion on the RMS Normalized AdaLN Module
5 Discussions and Future Works
In this paper, based on the successful low-bit training methods for large language models, we propose quantization-aware training (QAT) and efficient deployment methods for large-scale ternary DiT models. Competitive evaluation results on the ImageNet dataset (256$\times$256) demonstrate the feasibility of training a large ternary DiT from scratch while achieving results comparable to those of full-precision models. To the best of our knowledge, this work is also the first study concerning the quantization of DiT models.
While we believe this work provides valuable insights into the low-bit quantization of DiT models, it still has some limitations. Firstly, training ternary DiT is less stable and more time-consuming than full-precision networks. In our paper, although we discuss how to make the training more stable by adding norms, it still remains more time-consuming than training full-precision networks (Large-DiT-4.2B), which will lead to an increase in carbon dioxide emissions during model training in a broader context. Secondly, limited by computational resource constraints, we do not conduct ImageNet 512$\times$512 experiments, nor do we conduct experiments on the text-to-image generation task. However, we believe that the evaluation results on the ImageNet 256$\times$256 benchmark are already quite representative. The remaining tasks will be reserved for our future work. We hope that our work can reduce the deployment requirements of image generation models, and can inspire the community to join us in the future, fostering a broader advancement of this field of study.
Appendix A Appendix / supplemental material
A.1 Activation Distribution after Training
Here we conduct an analysis of the activation distribution during inference after model training, as a supplementary for the experiment and explanations in Sec. 3.3.
@@ A.1 Activation Distribution after Training
Following Sec. 4.3, we train the TerDiT-4.2B model with/without RMS Norm in the adaLN module for 50k steps. We then analyze the activation distribution of the ‘scale_mlp’ output in the adaLN module during inference, specifically at the first sampling step within the second ternary DiT block. For comparison, we also calculate the activation distribution of the original full-precision Large-DiT-4.2B model at the same layer. As shown in Fig. 7, training with RMS Norm can help limit the range of the activation values.
@@
Figure 7: Activation value analysis. We train the TerDiT-4.2B model with/without RMS Norm in the adaLN module for 50k steps and show the activation distribution of the ‘scale_mlp’ output (one output of the adaLN module) in the second ternary DiT block during inference (at the first sampling step). The activation distribution of the original full-precision model is also provided.
A.2 Comparison with Existing PTQ Methods
For a simple comparison with PTQ methods, we adopt the fake quant function for model weight quantization [^7] https://github.com/mit-han-lab/smoothquant/blob/main/smoothquant/fake_quant.py#L17 provided in SmoothQuant [45] and set n_bits=4. We do not carry out activation quantization. Same as TerDiT, we quantize all the linear layer weights in self-attention, feedforward, and MLP of the original DiT blocks. We then sample images with the 4-bit quantized model. In this case, both original DiT-XL/2 and Large-DiT-4.2B fail to generate images. As shown in Fig. 8. However, TerDiT can generate high-quality images with ternary precision. This highlights the significance of QAT for extremely low-bit DiT quantization.
Figure 8: DiT-XL/2 and Large-DiT-4.2B fail to generate images at 4-bit weight quantization.
A.3 Effectiveness of Learning Rate Reduction
In Sec. 4.1, we adopt a learning rate reduction after training for certain steps for more fine-grained parameter updates. Here we provide an ablation study on this learning rate reduction.
We choose the TerDiT-600M model for convenience. Following the setting of Sec. 4.1, we train a TerDiT-600M model with RMS Normalized adaLN module and 5e-4 learning rate for 1550k steps. We then continue training this model with 1e-4/5e-4 for another 200k steps.
Quantitative Results. We evaluate FID, sFID, Inception Score, Precision, and Recall of these two models. As can be seen in Tab. 3, the reduction in learning rate will lead to better evaluation results.
@@ A.3 Effectiveness of Learning Rate Reduction
Table 3: Effectiveness of learning rate reduction. Training with a reduced learning rate after 1550k steps for the TerDiT-600M model will lead to better evaluation results.
Qualitative Results. We also provide qualitative results comparison on TerDiT-600M with/without learning rate reduction in Fig. 9. For reference, the image generated by TerDiT-4.2B (fully trained) on the provided class label is shown in Fig. 10. The quality comparison highlights the importance of learning rate reduction in the later stages of training.
Figure 9: TerDiT-600M, class label [207, 360, 387, 974, 88, 979, 417, 279], cfg=4.
Figure 10: TerDiT-4.2B, class label [207, 360, 387, 974, 88, 979, 417, 279], cfg=4.
A.4 Qualitative Results Analysis for RMS Normalized AdaLN Module
In Sec. 4.3 we provide quantitative comparison results for models trained with/without RMS Norm in the adaLN module. Here we provide more qualitative results with models trained for 400k steps, together with the fully trained models.
@@ A.4 Qualitative Results Analysis for RMS Normalized AdaLN Module
Figure 11: Qualitative results comparison for 600M and 4.2B ternary DiT models with/without RMS Norm in the adaLN module. We choose class labels [972, 390, 417, 812, 118, 129, 447, 309] with cfg=4. Training the 4.2B model with the RMS Normalized adaLN module will lead to better qualitative generation results.
We compare TerDiT-600M and TerDiT-4.2B models trained with/without RMS Norm in the adaLN module. We sample images from the models trained for 400k steps in Sec. 4.3 and also show the results of the fully trained models. The comparisons are demonstrated in Fig. 11.
We can come to two conclusions:
  1. 1.
    For both the 600M and 4.2B TerDiT model, training with the RMS Normalized adaLN module will lead to better qualitative results.
  2. 2.
    Models with more parameters show more learning ability and can achieve better training results compared to models with fewer parameters.
These observations are consistent with the quantitative results provided in Sec. 4.3.
A.5 More Qualitative Results

### v2

@@ III-C QAT-specific Model Structure Improvement
The large-activation problem brought about by the ternary linear weights can be alleviated by applying a layer norm to the output of the ternary linear layer. We add an RMS Norm (similar to LLaMA) after the ternary linear layer and obtain the activation distribution in the left part of Fig. 4. In this case, the activation values are scaled to a reasonable range after passing the normalization layer and lead to more stable training behavior. The observation also aligns with [21], where a layer normalization function is applied before the activation quantization for each quantized linear layer.
TABLE I: Comparison between TerDiT and full-precision diffusion models on the ImageNet dataset.
@@
ImageNet 512×512 Benchmark
ADM [52] 1385 23.24 10.19 58.06 0.73 0.60
ADM-U [52] 1385 9.96 5.62 121.78 0.75 0.64
ADM-G [52] 1385 7.72 6.57 172.71 0.87 0.42
ADM-G, ADM-U [52] 1385 3.85 5.86 221.72 0.84 0.53
DiT-XL/2-G [7] 768 3.04 5.02 240.82 0.84 0.54
Large-DiT-4.2B-G [28] 472 2.52 5.01 303.70 0.82 0.57
@@
RMS Normalized AdaLN Module. We analyze the DiT model for QAT-specific model structure improvement based on the above insights. In a standard ViT transformer block, the layer norm is applied to every self-attention layer and feedforward layer. This is also the case with the self-attention layers and feedforward layers in the DiT block, which can help to properly scale the range of activations. However, the DiT block differs from traditional transformer blocks due to the presence of the adaLN module, as introduced in Sec. III-A. Notably, there is no layer normalization applied to this module. In the context of full-precision training, the absence of layer normalization does not have a significant impact. However, for ternary DiT networks, its absence can result in large dimension-wise scale and shift values in the adaLN (normalization) module, posing bad influences on model training. To mitigate this issue, we introduce an RMS Norm after the MLP layer of the adaLN module in each ternary DiT block:
$\text{adaLN}(c)=\text{RMS}(\text{MLP}(\text{SiLU}(c))),$ (5)
and the final model structure of TerDiT is illustrated in Fig. 3 (A). This minor modification can result in faster convergence speed and lower training loss, leading to better quantitative and qualitative evaluation results. To better showcase the effect, the actual activation distribution with/without the RMS Norm after model training is analyzed in Sec. IV-D, which shows similar distribution features.
Remark: What distinguishes TerDiT from classic low-bit quantization methods like BitNet b1.58 [22]?
In BitNet b1.58, a BitLinear layer is designed, which applies a layer norm to the input activation of the Linear layer. BitNet then replaces the Linear layers in LLaMA with BitLinear layers and removes the RMS Norm before the attention and SwiGLU layers. Similar approaches can be found in earlier works like Xnor-net [53], and Bi-real net [54] to construct a quantized layer. Different from applying layer norm to all the BitLinear layers of the model, TerDiT focuses on simple yet effective modifications to the model structure by simply adding an RMS Norm within the adaLN module of each DiT block. With fewer norms added, our method leads to faster training speed and better evaluation scores. Comparative experiments are introduced in Sec. IV-B.
III-D Deployment Scheme
Although there have been recent advancements in deploying ternary LLMs on CPUs [23], there are currently no effective open-source deployment solutions for ternary DiT models. This requires joint efforts from both the academic and engineering communities. To demonstrate the performance and deployment efficiency of ternary DiT and provide a reference for further advancing deployment on hardware, we deploy the trained networks with a 2-bit implementation. To be specific, we pack the ternary linear weights to INT8 values (4 ternary numbers into one INT8 number) with the pack_2bit_u8() function provided by [55]. During the inference process of the DiT model, we call the complementary unpack_2bit_u8() function on the fly to recover packed 2-bit numbers to FP values, then perform subsequent calculations. The addition of the unpacking operation will slow down the inference process, but we believe that with further research into model ternarization, more hardware support for inference speedup will be available.
IV Experiments
@@
Figure 5: Qualitative results analysis (256$\times$256). We compare images generated by DiT-XL/2 (A), Large-DiT-4.2B (B), and TerDiT-4.2B (C) with class labels [355, 980, 328, 33, 928, 862, 330, 944] and cfg=4. TerDiT-4.2B generates images of the same visual quality as two full-precision DiT models.
In this section, a set of experiments are designed to evaluate TerDiT. We compare TerDiT with full-precision diffusion models in Sec. IV-A, with other quantization methods (PTQ and QAT) in Sec. IV-B, carry out deployment efficiency comparison in Sec. IV-C, illustrate the effectiveness of the RMS Normalized adaLN module in Sec. IV-D, and carry out more ablation studies in Sec. IV-E. To further enrich the content of the paper, we show the generation results with high cfg values in Sec. IV-F and provide additional qualitative results in Sec. IV-G to visually demonstrate the capabilities of TerDiT. Our DiT implementation is based on the open-sourced code of Large-DiT-ImageNet [28] [^2] https://github.com/Alpha-VLLM/LLaMA2-Accessory/tree/main/Large-DiT-ImageNet. We conduct experiments on ternary DiT models with 600M (size of DiT-XL/2) and 4.2B [^3] The provided 3B model in the Large-DiT-ImageNet repository actually has 4.2B parameters. (size of Large-DiT-4.2B) parameters, respectively.
IV-A Comparison with Full-precision Models
We provide quantitative and qualitative comparison results between TerDiT and representative full-precision models in this subsection.
Experiment Setup. Following the evaluation setting of the original DiT paper [7], we train 600M and 4.2B ternary DiT models on the ImageNet dataset. We start from training the 256$\times$256 image resolution model (600M and 4.2B) and continue to train the 512$\times$512 resolution model (4.2B) based on the 256$\times$256 checkpoint. We compare TerDiT with a series of full-precision diffusion models, report FID [56], sFID [57], Inception Score, Precision, and Recall (50k generated images) following [52]. We also provide the total number of images (million) during the training stage as in [28] to offer further insights into the convergence speed of different models.
Training Details. For the 256$\times$256 resolution model, we train the 600M TerDiT model on 8 A100-80G GPUs for 1750k steps with batchsize set to 256, and the 4.2B model on 16 A100-80G GPUs for 1180k steps with batchsize set to 512. We set the initial learning rate as 5e-$4$The default learning rate of DiT is 1e-4. For TerDiT, we increase the initial learning rate to 5e-4 following [21] that a larger learning rate is needed for the faster convergence of low-bit QAT.. After 1550k steps of training for the 600M and 550k steps for the 4.2B model, we reduce the learning rate back to 1e-4 for more fine-grained parameter updates.
For 512$\times$512 resolution, we start from the 4.2B trained 256$\times$256 ternary model (trained with 604M images) and continue to train it on ImageNet with 512$\times$512 resolution and only 92M images. The learning rate is set to 1e-4.
TABLE II: Deployment efficiency comparison with cfg=1.5.
Resolution Checkpoint Size Max Memory Allocated Inference Time FID
DiT-XL/2-G 256 2.6GB 3128.42MB 15s 2.27
DiT-XL/2-G 512 2.6GB 3728.61MB 80s 3.04
Large-DiT-4.2B-G 256 16GB 17027.29MB 83s 2.10
Large-DiT-4.2B-G 512 16GB 18614.29MB 365s 2.52
[…]
TerDiT-4.2B-G 256 1.1GB 1919.67MB 97s 2.42
TerDiT-4.2B-G 512 1.1GB 3506.67MB 376s 2.81
Quantitative Results Analysis. The evaluation results are listed in Tab. III-C. TerDiT is a QAT scheme for DiT models, so among all the full-precision models, we pay special attention to DiT-XL/2 (675M) and Large-DiT-4.2B. (Large-DiT-4.2B and TerDiT-4.2B-G share the exact same architecture.)
In the 256$\times$256 case, without classifier-free guidance, TerDiT-4.2B achieves very similar testing results with DiT-XL/2 (with much fewer training images). With classifier-free guidance (cfg=1.5), TerDiT-4.2B-G outperforms LDM-G while bringing a very slight performance degradation compared with two full-precision DiT-structured models. Besides, TerDiT-4.2B-G achieves better evaluation results than TerDiT-600M-G, implying that models with more parameters can incur smaller performance degradation after quantization. In the more commonly used 512$\times$512 setting, TerDiT-4.2B-G outperforms DiT-XL/2-G in all aspects. It also surpasses Large-DiT-4.2B-G in terms of sFID and Precision. This result demonstrates the generative capabilities of TerDiT compared with full-precision models with the same architecture and parameter numbers.
Qualitative Results Analysis. To visually demonstrate the effectiveness of TerDiT, we also show some qualitative comparison results in Fig. 5 (256$\times$256), concerning TerDiT-4.2B, DiT-XL/2, and Large-DiT-4.2B. In terms of visual perception, there is no significant difference between the images generated by TerDiT and those by the full-precision models.
IV-B Extremely Low-bit Quantization Baselines
The topic of our paper is to explore an algorithm for the ternary (1.58-bit) quantization of extremely low-bit DiT models. In this section, we make a comparison between TerDiT and PTQ/QAT baselines.
PTQ Baselines. We choose PTQ algorithms Q-DiT [15] for DiT models and Q-Diffusion [11] for U-Net-based models and perform 2-bit weight quantization. We find they both fail to generate images, detailed examples are shown in Fig. 2.
QAT Baselines. [^5] We do not find code for replicating Q-DM [39] and they do not provide the ImageNet 256$\times$256/512$\times$512 results. We adapt BitNet b1.58 [21, 22] for the DiT model (256$\times$256, 600M parameters). We also use EfficientDM [37] for a more comprehensive analysis. For BitNet, we replace the linear layers in DiT (which correspond to TerDiT) with BitLinear and remove the norms before attention and SwiGLU layers. We then train BitNet using the same process as TerDiT and measure the FiD score (↓), which yields 6.60 for BitNet and 4.34 for TerDiT. An intuitive explanation for the phenomenon is that adding more norms can stabilize the training, but it may also slow down the convergence speed. Moreover, due to the additional norms, the training speed of BitNet drops to 0.9$\times$ that of TerDiT, while inference latency increases to 1.15$\times$. This demonstrates the efficiency of our proposed TerDiT. For EfficientDM, we apply 2-bit quantization and train the quantized model, but it fails to generate normal images, as detailed in Fig. 6.
IV-C Deployment Efficiency Comparison
Figure 6: Generation results of EfficientDM with 2-bit weight-only quantization.
@@ IV-C Deployment Efficiency Comparison
The improvement in deployment efficiency is the motivation of our proposed TerDiT scheme. In this subsection, we provide a comparison between TerDiT-600M/4.2B, DiT-XL/2, and Large-DiT-4.2B to discuss the actual deployment efficiency TerDiT can bring about. Tab. IV-A shows the checkpoint sizes of the four DiT models. We also record the memory usage and inference time of the total diffusion sample loop (step = 250) on a single A100-80G GPU.
@@
From the table, we can see that TerDiT greatly reduces checkpoint size and memory usage. The checkpoint size and memory usage of the 4.2B ternary DiT model are significantly smaller than those of Large-DiT-4.2B, even smaller than DiT-XL/2. This brings significant advantages to deploying the model on end devices (e.g., mobile phones, FPGA). However, the absence of open-source software deployment frameworks and efficient hardware support for ternary DiTs leads to slower inference speeds compared to their full-precision counterparts. Despite this, recent engineering advancements in deploying ternary LLMs on CPUs [23] suggest that the computational benefits of ternary-weight networks will become more apparent as software and hardware co-design continue to evolve.
@@
IV-D Ablation Study: RMS Normalized AdaLN Module
Experiment Setup. We train ternary DiT models with 600M and 4.2B parameters on the ImageNet [29] dataset in 256$\times$256 resolution. For each parameter size, we train two models, one with RMS Norm in the adaLN module and one without. We record the loss curves during training and measure the FID-50k score (cfg=1.5) every 100k training steps.
Training Details. For a fair comparison, we train all the ternary DiT models on 8 A100-80G GPUs with batchsize set to 256. The learning rate is set to 5e-4 throughout training.
Quantitative Results Analysis. The training loss [^6] The training process of diffusion models is not as ‘smooth’ as one might assume. To better illustrate the training dynamics, we employ exponential smoothing (with a smoothing factor of 0.995) during visualization. and evaluation scores are shown in Fig. 7 and Fig. 8 respectively. As can be seen, training with the RMS Normalized adaLN module will lead to faster convergence speed and lower FID scores. Another observation is that models with more parameters tend to achieve faster and better training compared to models with fewer parameters. This also, to some extent, reflects the scaling law of the ternary DiT model.
Qualitative Results Analysis. We compare the TerDiT-600M and TerDiT-4.2B models trained with/without RMS Norm in the adaLN module. We sample images from the models trained for 400k steps and also show the results of the fully trained models. The comparisons are demonstrated in Fig. 10. We can come to two conclusions:
1) For both the 600M and 4.2B TerDiT model, training with the RMS Normalized adaLN module will lead to better qualitative results.
2) Models with more parameters show more learning ability and can achieve better training results compared to models with fewer parameters.
@@ IV-D Ablation Study: RMS Normalized AdaLN Module
Activation Distribution after Training. We also conduct an analysis of the activation distribution during inference after model training as a supplementary for the experiment and explanations in Sec. III-C. We train the TerDiT-4.2B model (256$\times$256) with/without RMS Norm in the adaLN module for 50k steps and then analyze the activation distribution of the ‘scale_mlp’ output in the adaLN module during inference, specifically at the first sampling step within the second ternary DiT block. For comparison, we also calculate the activation distribution of the original full-precision Large-DiT-4.2B model at the same layer. As shown in Fig. 9, training with RMS Norm can help limit the range of the activation values.
@@
Figure 9: Activation value analysis. We train the TerDiT-4.2B model with/without RMS Norm in the adaLN module for 50k steps and show the activation distribution of the ‘scale_mlp’ output in the second ternary DiT block during inference (at the first sampling step). The activation distribution of the original full-precision model is also provided.
TABLE III: Effectiveness of learning rate reduction.
@@
Figure 10: Qualitative results comparison for 600M and 4.2B ternary DiT models with/without RMS Norm in the adaLN module. We choose class labels [972, 390, 417, 812, 118, 129, 447, 309] with cfg=4.
Figure 11: Qualitative comparison with/without lr reduction using TerDiT-600M. We choose class label [207, 360, 387, 974, 88, 979, 417, 279] and cfg=4.
IV-E Ablation Study: Learning Rate Reduction
In Sec. IV-A, we adopt a learning rate reduction after training for certain steps for more fine-grained parameter updates. Here, we provide an ablation study on this learning rate reduction.
We choose the TerDiT-600M model (256$\times$256) for convenience. Following the setting of Sec. IV-A, we train a TerDiT-600M model with the RMS Normalized adaLN module and 5e-4 learning rate for 1550k steps. We then continue training this model with 1e-4/5e-4 for another 200k steps.
Quantitative Results. We evaluate FID, sFID, Inception Score, Precision, and Recall of these two models. As can be seen in Tab. III, the reduction in the learning rate will lead to better evaluation results.
Qualitative Results. We also provide qualitative results comparison on TerDiT-600M (256$\times$256) with/without learning rate reduction in Fig. 11.
The quality comparison highlights the importance of learning rate reduction in the later stages of training.
IV-F Image Quality with High CFG Values
In our experiments above, we use a cfg scale of 4 for qualitative evaluations. Setting the cfg scale too high can result in images that are less creative and of lower quality. In this section, we test the robustness of TerDiT when handling high cfg scale. We provide sample images with cfg=4 and cfg=10 in Fig. 12. Although there is slight distortion, quantization has not severely impacted the generation quality at high cfg values, demonstrating the robustness of TerDiT to extreme cases.
IV-G More Qualitative TerDiT Results
In this subsection, we provide more generation results of TerDiT with 256$\times$256 resolution in Fig. 13, with 512$\times$512 resolution in Fig. 14. TerDiT is capable of generating images with good visual quality.